What are the values of the following expressions?

  • ∑_(i=3)^6▒ (i^2+6)
  • ∏_(j=1)^5▒ (j/2)
  • 4!
  • 0!
  • (6¦2)
  • C(6,2)
1
Expert's answer
2021-12-01T18:23:01-0500

(i)


"\displaystyle\sum_{i=3}^6(i^2+6)=3^2+6+4^2+6+5^2+6+6^2+6"

"=110"

(ii)


"\displaystyle\prod_{j=1}^5(i/2)=\dfrac{1(2)(3)(4)(5)}{2(2)(2)(2)(2)}=\dfrac{15}{4}"

(iii)


"4!=1(2)(3)(4)=24"

(iv)


"0!=1"

(v)


"P(6, 2)=\dfrac{6!}{(6-2)!}=6(5)=30"

(vi)


"C(6, 2)=\dfrac{6!}{2!(6-2)!}=\dfrac{6(5)}{1(2)}=15"